A 2700-kg Ferris wheel accelerates from rest to an angular speed of 0.16 rad/s in 14 s . Approximate the Ferris wheel as a circular disk with a radius of 23 m . What is the net torque on the wheel?

Net Torque = _____ m*N
Answer #1

Solution)

Mass,m=2700 kg

Angular speed, v=0.16 rad/s

Time, t=14 s

Radius,r=23 m

Here, Moment of Inertia, I= 0.5 mr^2= 0.5*2700*23^2

=714150 kg-m^2

Also, angular acceleration, alpha=0.16/14= 0.0114 rad/s^2

Now, we know

Torque, T= I*alpha=714150*0.0114= 8161.7 N=8161700 mN (Ans)

==============